OpenStudy (anonymous):

What is the general form of the equation of a circle with center at (a, b) and radius of length m? A) x2 + y2 − 2ax − 2by + (a2 + b2 − m2) = 0 B) x2 + y2 + 2ax + 2by + (a2 + b2 − m2) = 0 C) x2 + y2 − 2ax − 2by + (a + b − m2) = 0 D) x2 + y2 + 2ax + 2by + a2 + b2 = -m2

OpenStudy (anonymous):

\[ x^2 + y^2 − 2ax − 2by + (a^2 + b^2 − m^2) = 0 \\ x^2 - 2 ax + a^2 + y^2- 2 by +b^2 = m^2\\ (x-a)^2 + (y-b)^2=m^2 \] Is this the right equation?

OpenStudy (anonymous):

IS IT B?

OpenStudy (anonymous):

No, try again

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

Compare my first line to the choices

OpenStudy (anonymous):

a?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thank you